OpenStudy (anonymous):

Find the equation of the line tangent to the function f(x)=5x^2 at x =10. *not toooo sure how to do this, but would I just be subbing in 10 for x? f(1)=5(10)^2 =5(100) =500? or is it something else? please explain! thanks :)

OpenStudy (anonymous):

Find the slope first. Do you know how to do that?

OpenStudy (anonymous):

y2-y1/x2-x1 ? or would i be subbing the x value into a y=mx+b format? haha not sure in this scenario!

OpenStudy (anonymous):

What class are you in right now?

OpenStudy (anonymous):

calculus :)

OpenStudy (anonymous):

Use a calculus concept you learned not too long ago :)

OpenStudy (anonymous):

okay.. ermm am i using this? lim f(a+h)-f(a) h->0 ___________ h ? :/

ganeshie8 (ganeshie8):

it works, or u cud use the formula directly : slope = derivative derivative of \(\large x^n = nx^{n-1}\)

OpenStudy (anonymous):

ohh okay :) how would i use that here? would n=5 and x=10?

ganeshie8 (ganeshie8):

slope of \(x^2\) at any point (x, y) = derivative of \(\large x^2 = 2x^{2-1} = 2x\)

ganeshie8 (ganeshie8):

since 5 is a constant, u can pull it out. \(\large \frac{d}{dx}(5x^2) = 5*\frac{d}{dx}(x^2) = 5*2x = 10x\)

ganeshie8 (ganeshie8):

familiar wid this d/dx notation ?

OpenStudy (anonymous):

ahh okay, yes looks familiar :)

ganeshie8 (ganeshie8):

good :) so, slope of 5x^2 at any point (x, y) is 10x.

OpenStudy (anonymous):

so then now you just end up with 10(10)=100? and also just making sure... d=derivative right?

ganeshie8 (ganeshie8):

yes, and yes. so u have these at hand : slope, m = 100 point = (10, ??) u can write the equation of line ?

OpenStudy (anonymous):

ah okay :) umm was that from subbing in earlier? so y=500 y=mx+b slope = 2(10)=20 500=20(10)+b 500=200+b subtract 200 from both sides 300=b y=20x+300 ?

ganeshie8 (ganeshie8):

slope = 100 right ? lol how come it bacame 20 all off sudden :o

OpenStudy (anonymous):

haha ooops! ermm so is it 500=100(10)+b -500 =b? y=100x-500?

ganeshie8 (ganeshie8):

looks good !!

OpenStudy (anonymous):

okay awesome!! thank you :)

ganeshie8 (ganeshie8):

np... u wlc :)